OpenStudy (kweb2017):

Find the slope and the y-intercept of the equation y − 3(x − 1) = 0. slope= -3 and y-intercept -3 slope= 3 and y-intercept 3 slope= -3 and y-intercept 3 slope= 3 and y-intercept -3

OpenStudy (anonymous):

arrange it into\[y=mx+c\]

OpenStudy (anonymous):

where\[y=vertical~axis\]\[x=horizontal~axis\]\[m=slope\]\[c=y-intercept\]

OpenStudy (anonymous):

\[y\color{red}{-3(x-1)}=0\]First expand the left side

OpenStudy (anonymous):

\[y-3x+3=0\]Then,add both sides by 3x

OpenStudy (anonymous):

\[y-3x+\color{red}{3x}+3=\color{red}{3x}\]

OpenStudy (anonymous):

\[y+3=3x\]

OpenStudy (anonymous):

Then,subtract 3 on both sides

OpenStudy (anonymous):

\[y+3\color{red}{-3}=3x\color{red}{-3}\]

OpenStudy (anonymous):

\[y=3x-3\]

OpenStudy (anonymous):

now try comparing with\[y=mx+c\]where\[y=vertical~axis\]\[x=horizontal~axis\]\[m=slope\]\[c=y-intercept\]

OpenStudy (anonymous):

to find the slope and the y-intercept

OpenStudy (anonymous):

@kweb2017